UT 26 Feb 2019

OSA: D.G.Huerta

ISp: J. Morris

Mtn Manager: M Wagner

ISA: B. Rothberg (Tucson)

AO: S. Walsh (Tucson)

Observers: B. Rothberg, O. Kuhn, G. Rahmer, J. Hill (all Tuson)

ARGOS LOG: https://docs.google.com/document/d/1wXB5oGJZQeJZyvcKG0KmcRuf4j8PY46m9e3SG0pbUuQ/edit?usp=sharing


UT 00:26 - Error with LUCI-1 MCU - camera wheel error occurred, brought up Camera GUI and LUCI-1 MCU on Eng Panel - unable to connect. J. Morris investigating - resembles IT 7049 (but that was for LUCI-20

UT 01:14 - thick clouds overhead, not opening - concurrent problem, ARGOS laser system com error

UT 01:16 - on phone with I. Sola to diagnose whether the issue is the same as before

UT 01:25 - LUCI-1 HIRAMO being stopped/restarted, also stopping and starting LUCI-2 HIRAMO

UT 01:42 - Telescope being locked out - J. Morris going to check using LBT-LUCIFER-MAN-011 v4.1 - MOCON 1 appears to be the issue as I. Sola cannot even communicate with the system (IT 7049 the issue was MOCON 2, we could telnet in and the responses indicated it was MOCON 2)

UT 01:48 - Both LUCI AGWs not connecting to GCS. - could ARGOS, LUCI-1, and AGW all be a network issue?

UT 02:04 - no power lights on any MOCON boards on LUCI-1 (from J. Morris on the scope) - power switch is in ON position - this sounds like a broken fuse (PG 36 of 011 Manual)

UT 02:12 - contacted J. Riedl - some possibiities on lack of power - will contact J. Morris directly - still working on BCU/ARGOS issue

UT 02:47 - still closed - LUCI-1 offline while J. Morris and J. Riedl work on the MCU issue, SX ARGOS com issues being worked on, LUCI AGW1/2 problems

UT 02:52 - O. Kuhn working on MODS scripts for pointing

UT 03:15 - still closed - looks like LUCI-1 is not available for tonight ... will open an IT

UT 03:59 - IT 7583 - LUCI-1 MCU issue, D.G. Huerta and S. Hooper still working on BCU SX issue. The goal is to close the loop on the SX side even though we can't take any data with LUCI-1.

UT 04:05 - clouds....S. Hooper implies a damaged board affecting the SX BCU

UT 04:35 - still closed with clouds - seems like SX BCU issue can't be resolved - working on AGW com issue.

UT 05:14 - still working on AGW issues

UT 05:53 - still closed - working on AGW issues

UT 06:02 - plan is to start binocular - turn off instrument and detector on LUCI-1.

UT 06:10 - opening enclosure - powering on MODS-Bino just in case...

UT 06:59 - laser system on DX powered up - we will only be using the DX side tonight - unable to perform work with SX.

UT 07:00 - pointing check

UT 07:33 - haivng difficulties collimating - D. Miller working with D.G.Huerta to improve. DIMM ~1.4" (WFS showing 2.8")

UT 07:36 - sending ARGOS L2 Collimation script

UT 07:40 - collimation script complete - need to engage the laserization


UT 07:53 - still diagnosing laser system - GUIs don't seem to be refreshing...not enough power to lasers?

UT 08:10 - G. Rahmer working on diagnosing laser system issue.

UT 08:31 - class 1 laser system was deactivated/disabled - it was originally enabled but key was turned which powered it down

UT 08:37 - laser issue sorted

UT 08:44 - moving to M53 - its rising and M67 is at EL < 45

UT 08:54 - kept sending a Track preset even though the OT showed the setup as Active, diff'd the file with the successful M67 Collimation script and the guidestar info was missing, as was the Active preset...have no idea why.

UT 08:55 - dealing with another issue relating to the recovery of DX ARGOS - getting pop up dialogue box "ARGOS slopes not received by ASM" - restarted arbitrator - same error

UT 09:07 - D. Miller stopping and restarting AdSec (DX) to bring everything up cleanly

UT 09:21 - laser is propagating

UT 09:24 - sent observing script

UT 09:34 - pointing check, resent collimation preset with different ARGOS GS

UT 09:37 - found the AO Ref star - pointing was the issue

UT 09:43 - AO Ref Star should be R=10.3

UT 09:45 - in a long satellite closure, but sending to a new star R=13.4

UT 09:50 - see ripples - stars are elongated

UT 09:55 - resending preset

UT 10:01 - beam sharpening done by S. Walsh - resent preset

(L2 0001-0003 - loops opened or other issues)

UT 10:05 - loops opened towards end of first exposure (L2 0004)

UT 10:14 - S. Walsh looking for lasers....

UT 10:26 - still adjusting the lasers (they've been found)

UT 10:30 - resending preset

L2 0005 - loops stayed close (0,0)

L2 0006 - -20 +20 dither

UT 10:39 - opening loops

UT 10:55 - iterating with W. Gaessler on isssue with laser - open loop images look OK - closing loops with ARGOS introduces the issue

UT 11:17 - starting up lasers on SX to compare with DX

UT 11:31 - W. Gaessler turned off vibration compensation on DX - did get worse

UT 12:11 - still working on comparison between SX and DX launch mirrors and focus

UT 12:51 - attempting binocular pointing - getting mulitple co-pointing errors

-sent bino col preset and bino script (with SX/LUCI-1 insturment and readout deactivated) and lasers look better on DX

L2 0007-0009 - 2.51 sec indvidual exposures

L2 0010 - 1 x 52 x 2.51 cube with data

UT 13:18 - beam sharpening

UT 13:21 - sky background too high for NGS ref star

UT 13:22 - putting LUCI-2 into safe config - LUCI-1 left as it (mostly safe)

UT 13:23 - closing enclosure - done for the night - putting MODS to sleep

-- BarryRothberg - 26 Feb 2019
Topic revision: r1 - 26 Feb 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback